# for Multimodal Understanding

### v1

@@ 5.3 Probability Distribution over Text Tokens Analysis
To better understand the behavior of Align, we examine the probability distribution, $\mathbf{P}_{\text{vocab}}$ in Eq (1), over the LLM’s text vocabulary generated from visual features. Specifically, we process 100 document images through the vision encoder and Align, then average the resulting probability distributions across all image patches. The final distribution is shown in Figure  3. As illustrated, the distribution is dense (rather than sparse), with the highest probability assigned to a single token being 0.0118. This can be explained by the vision feature space being continuous and of much higher cardinality than the discrete text space. Indeed, while the LLM has 128K distinct vocabulary tokens, an image patch (e.g., 14×14 pixels) contains continuous, high-dimensional information that cannot be effectively mapped to a single or a few discrete tokens.
Figure 3: Probability distribution over the LLM text tokens, showing dense probabilities and higher values for tokens associated with white space in document images.
Furthermore, we observe that tokens on the left side of the distribution in Figure  3 have higher probabilities than the rest. Upon investigation, we found that these tokens correspond to patches that are predominantly white – a common feature in document images. Further analysis of the associated text tokens reveals that they predominantly consist of punctuation marks, as illustrated further in Appendix A.2. This suggests that the model repurposes punctuation marks to represent whitespaces. This may be attributed to the fact that both punctuation and whitespaces act as structural cues and separators. Other possibilities include whitespaces being rarely directly-required to perform a task, and LLMs may pay less specific attention to common tokens such as punctuation.
@@
5.4 Pixel-Level Tasks Analysis
To rigorously evaluate the ability of vision-language models to integrate fine-grained visual and textual pixel-level cues, we test our model on the VCR benchmark (Zhang et al., 2024), which requires the model to recover partially occluded texts with pixel-level hints from the revealed parts of the text. This task challenges VLM’s alignment of text and image in extreme situations. Current state-of-the-art models like GPT-4V (OpenAI et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024), and Llama-3.2 (Dubey et al., 2024) significantly underperform humans on hard VCR task due to their inability to process subtle pixel-level cues in occluded text regions. These models frequently discard critical visual tokens during image tokenization on semantic priors, overlooking the interplay between partial character strokes and contextual visual scenes. To evaluate performance on VCR, we modify our Stage 3 SFT dataset composition by replacing the exclusive use of DocDownstream with a 5:1 blended ratio of DocDownstream and VCR training data. This adjustment enables direct evaluation of our architecture Align’s ability to leverage pixel-level character cues.
Figures 5(c) and 5(d) show examples where AlignVLM fails on the VCR task. These carefully picked instances show that our method mistakes names of landmarks with common words when the two are very similar. As seen in the examples, AlignVLM mistakes “Llanengan” for “Llanongan” and “Gorden” for “Garden”. In both instances, the pairs differ by one character, indicating perhaps that AlignVLM tends to align vision representations to more common tokens in the vocabulary. One approach that would potentially mitigate such errors would be to train AlignVLM with more contextually-relevant data.
5.5 Robustness to Noise Analysis
Model Without Noise With Noise Drop ($\Delta$)
Llama-3.2-3B-MLP 53.06 27.52 $\downarrow 25.54$
Llama-3.2-3B-Align (ours) 58.81 57.14 $\downarrow\textbf{1.67}$
As shown in Table 3, our Align connector demonstrates high robustness to noise, with only a 1.67% average drop in performance. In contrast, the widely adopted MLP connector suffers a significant performance degradation of 25.54%, highlighting its vulnerability to noisy inputs. These empirical results support our hypothesis that leveraging the knowledge encoded in the LLM’s text embeddings and constraining the visual features within the convex hull of the text latent space act as a regularization mechanism, reducing the model’s sensitivity to noisy visual features.
6 Conclusion
We introduce Align, a novel connector designed to align vision and language latent spaces in vision-language models (VLMs), specifically enhancing multimodal document understanding. By improving cross-modal alignment and minimizing noisy embeddings, our models, AlignVLM, which leverage Align, achieve state-of-the-art performance across diverse document understanding tasks. This includes outperforming base VLMs trained on the same datasets and open-source instruct models trained on undisclosed data. Extensive experiments and ablations validate the robustness and effectiveness of Align compared to existing connector designs, establishing it as a significant contribution to vision-language modeling. Future work will explore training on more diverse instruction-tuning datasets to generalize beyond document understanding to broader domains.
Impact Statement
This work contributes to the multimodal AI research community by introducing a novel approach for fusing vision and language modalities within large language models. By leveraging our framework in the context of generative models, we enable more effective integration of visual and textual information, enhancing the generative models’ ability to generate free-form text for multimodal tasks. However, like all generative models, our approach is subject to potential biases and hallucinations—challenges inherent to large language models that must be carefully considered in deployment. Since these issues are not unique to our approach, we do not highlight any specific concerns here.
Appendix A Appendix
A.1 Experimental Setup
We provide detailed hyperparameters of our experiments in Table 4.
LLM Backbone Llama 3.2-1B Llama 3.2-3B Llama 3.1-8B Stage-1 Stage-2 Stage-3 Stage-1 Stage-2 Stage-3 Stage-1 Stage-2 Stage-3 Trainable Parameters Full Model Full Model LLM & Align Full Model Full Model LLM & Align Full Model Full Model LLM & Align Batch Size 512 512 512 512 256 256 512 256 256 Text Max Length 1024 2048 2048 1024 2048 2048 1024 2048 2048 Epochs 1 1 5 1 1 5 1 1 5 Learning Rate $1\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
A.2 Vision-to-Text
In this experiment, we analyze how Align maps visual features to the LLM’s text tokens. To do so, we manually curate a small dataset of image crops, each containing either a single word or a small set of visual text elements. Unlike the processing of high-resolution images described earlier (Section 3.1), these image crops are not divided into tiles. Instead, the backbone image encoder processes each crop as a single tile, producing $14\times 14$ features from the input image. The resulting features pass through the Softmax operation (Equation 1), yielding a probability distribution over the LLM’s text tokens for each feature (region). We examine the decoded text tokens from specific image regions to better understand how visual features are mapped to textual representations.
As shown in Figure 6, white regions in the images tend to assign higher probabilities to punctuation tokens, such as commas or periods. Since punctuation structures written text, while white space separates document components like paragraphs, tables, and sections, Align appears to leverage these implicit patterns to align visual structures with semantically meaningful representations in the LLM’s embedding space.
Figure 6: Mapping Visual-to-Text tokens. The left column shows the visual input to the model. In contrast, the right column visualizes the decoded tokens on a 14×14 grid, displaying the top k=2 tokens corresponding to the most likely LLM tokens predicted for the respective visual feature in each cell.
A.3 Case Studies
@@
In this section, we provide case studies for the experiments in Section 5.1. Specifically, we provide examples of our Llama-3.2-3B-Align, and its counterpart model with alternative connectors Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis on three different datasets: KLC (Stanisławek et al., 2021), DocVQA (Mathew et al., 2021b), and TextVQA (Singh et al., 2019). The examples are shown in Figure 7, 8, and 9.
@@
Figure 7: Case Study for Connector Comparison on the KLC dataset (Stanisławek et al., 2021). We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.
@@
Figure 8: Case Study for Connector Comparison on the DocVQA dataset (Mathew et al., 2021b). We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.
@@
Figure 9: Case Study for Connector Comparison on the TextVQA dataset (Singh et al., 2019). We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.

### v2

@@ 5.3 Probability Distribution over Text Tokens Analysis
To better understand the behavior of Align, we examine the probability distribution, $\mathbf{P}_{\text{vocab}}$ in Eq (1), over the LLM’s text vocabulary generated from visual features. Specifically, we process 100 document images through the vision encoder and Align, then average the resulting probability distributions across all image patches. The final distribution is shown in Figure  4. As illustrated, the distribution is dense (rather than sparse), with the highest probability assigned to a single token being 0.0118. This can be explained by the vision feature space being continuous and of much higher cardinality than the discrete text space. Indeed, while the LLM has 128K distinct vocabulary tokens, an image patch (e.g., 14×14 pixels) contains continuous, high-dimensional information that cannot be effectively mapped to a single or a few discrete tokens.
Table 4: Performance comparison when evaluating Align with the full text embedding vocabulary (128K) versus the reduced subset of 3.4K high-probability embeddings. The results show negligible performance degradation, indicating that Align relies primarily on a small subset of embeddings.
Llama-3.2-3B-Align (Full Embeddings) 79.63 44.53 63.49 35.25 38.59 78.51 71.88 57.38 60.10 58.81
Llama-3.2-3B-Align (3.4K Embeddings) 79.40 44.13 63.64 35.02 38.26 78.83 71.72 57.48 59.80 58.69
We conducted a deeper analysis of the token probability distributions produced by the Align connector. Our observations show that Align consistently assigns high probabilities to approximately 3.4K tokens from the entire vocabulary, while the remaining tokens receive negligible probabilities (below $10^{-6}$). To better understand this behavior, we applied Principal Component Analysis (PCA) to reduce the dimensionality of the embeddings and visualized them in a two-dimensional space, as shown in Figure 4. The visualization reveals that these 3.4K tokens densely and comprehensively span the latent space of the LLM’s text embeddings. To validate this finding, we conducted additional evaluation experiments in which we retained only these 3.4K high-probability embeddings in the Align connector, entirely removing the rest during evaluation. As shown in Table 4, the performance difference compared to using the full embedding set (128K) was negligible. This confirms that Align effectively leverages and combines a compact subset of embeddings to map visual features into semantically meaningful regions within the LLM’s latent text space. Moreover, this suggests that Align can be further optimized through targeted embedding pruning to improve computational efficiency without sacrificing performance.
Figure 3: Probability distribution over LLM tokens, highlighting dense probabilities for whitespace tokens.
Figure 4: PCA of Align Embeddings: The principal components of the most influential embeddings in the Align Connector span most of the feature space represented by all embeddings.
5.4 Robustness to Noise Analysis
Model Without Noise With Noise Drop ($\Delta$)
Llama-3.2-3B-MLP 53.06 27.52 $\downarrow 25.54$
Llama-3.2-3B-Align (ours) 58.81 57.14 $\downarrow\textbf{1.67}$
As shown in Table 5, our Align connector demonstrates high robustness to noise, with only a 1.67% average drop in performance. In contrast, the widely adopted MLP connector suffers a significant performance degradation of 25.54%, highlighting its vulnerability to noisy inputs. Furthermore, we measured the average cosine distance between the original and noise-perturbed visual embeddings using both the Align and MLP connectors. Align showed significantly lower distances (0.0036) than MLP (0.3938), further validating its robustness to noise. These empirical results support our hypothesis that leveraging the knowledge encoded in the LLM’s text embeddings and constraining the visual features within the convex hull of the text latent space act as a regularization mechanism, reducing the model’s sensitivity to noisy visual features.
6 Conclusion
We introduce Align, a novel connector designed to align vision and language latent spaces in vision-language models (VLMs), specifically enhancing multimodal document understanding. By improving cross-modal alignment and minimizing noisy embeddings, our models, AlignVLM, which leverage Align, achieve state-of-the-art performance across diverse document understanding tasks. This includes outperforming base VLMs trained on the same datasets and achieving competitive performance with open-source instruct models trained on undisclosed data. Extensive experiments and ablations validate the robustness and effectiveness of Align compared to existing connector designs, establishing it as a significant contribution to vision-language modeling. Future work will explore training on more diverse instruction-tuning datasets to generalize to broader domains.
Appendix A Appendix
A.1 Experimental Setup
We provide detailed hyperparameters of our experiments in Table 6.
LLM Backbone Llama 3.2-1B Llama 3.2-3B Llama 3.1-8B Stage-1 Stage-2 Stage-3 Stage-1 Stage-2 Stage-3 Stage-1 Stage-2 Stage-3 Trainable Parameters Full Model Full Model LLM & Connector Full Model Full Model LLM & Connector Full Model Full Model LLM & Connector Batch Size 512 512 512 512 256 256 512 256 256 Text Max Length 1024 2048 2048 1024 2048 2048 1024 2048 2048 Epochs 1 1 5 1 1 5 1 1 5 Learning Rate $1\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-5}$ $1\times 10^{-5}$
A.2 Runtime Comparison Between Connectors
One caveat in the Align connector is that it includes an additional LM head layer, which slightly increases the total number of parameters. However, this addition has a negligible impact on runtime efficiency due to its simple structure. It only introduces a few matrix multiplication operations (as shown in Equations 1 and 2) instead of stacking many complex layers that require sequential processing, as in deep fusion methods.
To empirically validate this claim, we benchmarked the runtime and memory usage of models equipped with different connector types (MLP, Align, Ovis, and Perceiver), following the same experimental setup as in Table 2. As shown in Table 7, the results demonstrate that although the Align connector delivers notably superior performance (see Table 2), the variations in inference speed and GPU memory usage among the connectors remain minimal.
Table 7: Runtime and memory comparison between different connector designs. The results show that Align introduces negligible computational overhead compared to other connectors.
Model Samples Avg Time (s) Tokens/sec GPU Memory (GB)
Llama-3.2-3B-MLP 2500 0.161 118.3 10.9
Llama-3.2-3B-Perceiver 2500 0.140 135.1 10.9
Llama-3.2-3B-Ovis 2500 0.155 122.5 10.8
Llama-3.2-3B-Align 2500 0.165 115.4 10.9
Overall, the empirical evidence confirms that the Align connector achieves an effective balance between computational efficiency and performance. It introduces only a negligible increase in runtime and memory usage while providing substantial gains in overall accuracy.
A.3 Pixel-Level Tasks Analysis
To rigorously evaluate the ability of vision-language models to integrate fine-grained visual and textual pixel-level cues, we test our model on the VCR benchmark [Zhang et al., 2024], which requires the model to recover partially occluded texts with pixel-level hints from the revealed parts of the text. This task challenges VLM’s alignment of text and image in extreme situations. Current state-of-the-art models like GPT-4V OpenAI et al. [2023], Claude 3.5 Sonnet Anthropic [2024], and Llama-3.2 Dubey et al. [2024] significantly underperform humans on hard VCR task due to their inability to process subtle pixel-level cues in occluded text regions. These models frequently discard critical visual tokens during image tokenization on semantic priors, overlooking the interplay between partial character strokes and contextual visual scenes. To evaluate performance on VCR, we modify our Stage 3 SFT dataset composition by replacing the exclusive use of DocDownstream with a 5:1 blended ratio of DocDownstream and VCR training data. This adjustment enables direct evaluation of our architecture Align’s ability to leverage pixel-level character cues.
Figures 6(c) and 6(d) show examples where AlignVLM fails on the VCR task. These carefully picked instances show that our method mistakes names of landmarks with common words when the two are very similar. As seen in the examples, AlignVLM mistakes “Llanengan" for “Llanongan" and “Gorden" for “Garden”. In both instances, the pairs differ by one character, indicating perhaps that AlignVLM tends to align vision representations to more common tokens in the vocabulary. One approach that would potentially mitigate such errors would be to train AlignVLM with more contextually-relevant data.
@@ A.3 Pixel-Level Tasks Analysis
A.4 Case Studies
@@
In this section, we provide case studies for the experiments in Section 5.1. Specifically, we provide examples of our Llama-3.2-3B-Align, and its counterpart model with alternative connectors Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis on three different datasets: KLC [Stanisławek et al., 2021], DocVQA [Mathew et al., 2021b], and TextVQA [Singh et al., 2019]. The examples are shown in Figure 7, 8, and 9.
@@
Figure 7: Case Study for Connector Comparison on the KLC dataset [Stanisławek et al., 2021]. We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.
@@
Figure 8: Case Study for Connector Comparison on the DocVQA dataset [Mathew et al., 2021b]. We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.
@@
Figure 9: Case Study for Connector Comparison on the TextVQA dataset [Singh et al., 2019]. We show four qualitative examples (including two correct and two incorrect examples) comparing Llama-3.2-3B-Align to the same architecture with different connectors, Llama-3.2-3B-MLP and Llama-3.2-3B-Ovis. “GT” denotes the ground truth.